# Magnetisation moment of a bounded 3D sample: asymptotic recovery from planar measurements on a large disk

### v4

@@ Abstract.
We consider the problem of reconstruction of the net moment (vector of the overall magnetisation) of a sample from partial data of the magnetic field. Namely, motivated by a concrete experimental set-up, we deal with a situation when the magnetic field is measured on a portion of the plane in vicinity of the sample and only one (normal to the plane) component of the field is available. Under assumption that the measurement area is a sufficiently large disk (lying in a horizontal plane above the sample), we obtain a set of estimates for the components of the net moment vector with the accuracy which improves asymptotically with the increase of the measurement disk radius. Compared to our previous preliminary results, the asymptotic formulas are now rigorously justified and higher-order estimates are derived. Moreover, the presented approach, based on an appropriate splitting in the Fourier domain and estimates of oscillatory integrals (involving both small and large parameters), elucidates the derivation of asymptotic estimates of an arbitrary order, a possibility that was previously unclear. The obtained results are illustrated numerically and their robustness with respect to noise is discussed. The proposed methodology should be applicable to other magnetic and gravimetric problems with planar measurements.
@@ 1. Introduction
In the present work, we are concerned with recovery of the overall magnetisation (the so-called net moment) of a sample rather than dealing with reconstruction of the entire magnetisation distribution. While both inverse problems are known to be ill-posed (as an inverse source problem for elliptic partial-differential equations) due to the lack of continuous dependence of their solution on the input data (magnetic field measurements), the problem of reconstruction of full magnetisation distribution additionally lacks uniqueness of the solution in view of presence of invisible (or “silent”) sources, i.e., magnetisations that do not produce magnetic field, see [4]. However, as it was shown in [2] for planar (thin-plate) magnetisation distributions, and as follows from the general Helmholtz decomposition [7], compactly supported invisible sources do not contribute to the net moment. This statement fixes the non-uniqueness issue and makes the problem of net moment recovery a feasible task. In theory, this problem is even solvable in a closed form when measurements are available on the entire plane above the sample. In reality, however, the measurements are very limited and corrupted by the presence of noise which dominates the signal in distant regions. Therefore, we arrive at the problem of estimating the net moment of a sample from a magnetic field component available on some portion of the plane in proximity of the sample. We shall assume that this portion of the plane is sufficiently large, for otherwise the instability of the problem, due to the already mentioned inherent ill-posedness, will be even more severe. On the other hand, this assumption on the large size of the measurement area allows us to obtain explicit ready-to-use formulas.
@@
(1.1) $B_{3}\left(\vec{x}\right)=\dfrac{\partial}{\partial x_{3}}\iiint_{Q}\dfrac{1}{4\pi\left|\vec{x}-\vec{t}\right|}\nabla\cdot\vec{\mathcal{M}}\left(\vec{t}\right)\mathrm{d}^{3}t,\hskip 10.00002pt\vec{x}\in\mathbb{R}^{3}\backslash\overline{Q}.$
@@
This is a constant $\mathbb{R}^{3}$ vector equal to the overall magnetisation of the sample which can be determined uniquely, unlike the full magnetisation distribution $\vec{\mathcal{M}}$ which essentially enters $B_{3}$ through $\nabla\cdot\vec{\mathcal{M}}$ and hence can be changed by any divergence-free vector field without altering $B_{3}$, see (1.1) and discussions in [4]. In what follows, we may intechangeably use slightly different names for $\vec{m}$: net moment, net magnetisation, zeroth-order algebraic moment of a magnetisation distribution. When $\vec{\mathcal{M}}$ is not a regular function but a distribution, the integral in (1.3) should be understood componentwise as the duality pairing of a compactly supported distribution with the constant function $1$ on $\mathbb{R}^{3}$: $m_{j}=\left\langle\mathcal{M}_{j},1\right\rangle$, $j\in\left\{1,2,3\right\}$.
@@
[…]
@@
We stress that our analytical expressions for the asymptotic estimates are original, with only partial analogs in other works when translated into the current context. According to [20], for extended finite magnetic sources, net moment estimates in terms of the measured field have first appeared in Helbig’s work [12] and later rediscovered by Clark and Schmidt, see, e.g., [23]. When truncated to a finite measurement area, these estimates correspond to our first-order asymptotic formulas for tangential components of net moment, but without a mention of the order of the asymptotic convergence, let alone a possibility of obtaining their more precise higher-order counterparts. Note that, in our context, since only $B_{3}$ component of the field is measured, those previous estimates do not provide means to estimate the normal net moment component $m_{3}$. Indeed, an estimate of $m_{3}$ was provided only in terms of either $B_{1}$ or $B_{2}$ components. When measurements of these two field components are not available, one can potentially argue that it is still possible to use the Helbig’s formulas after reconstructing the missing components of the field from the knowledge of $B_{3}$ alone using, for example, [11] and [8, Ch. 12]. However, this is possible only to accomplish in a stable way only if $B_{3}$ is known on the entire plane. In view of this, we believe that an explicit estimate of $m_{3}$ from knowledge of $B_{3}$ alone on a finite region for the first time was proposed only in [21, 5, 3]. It should be noted that results in [3] are also applicable for measurement regions different from disk, namely, for those of rectangular and diamond shapes. We take advantage of this comparison of results to mention and to correct a couple of issues in [5]. First of all, there is a typo in the $m_{3}$ formula obvious when compared to (2.4) or [21, Thm 3.1.1]: in eq. [5, Eq. (5)], the factor $A$ should be in the numerator rather than the denominator. Moreover, the numerical illustration of the $m_{3}$ estimate, in case of noisy data, showed a divergent behavior with the growth of the measurement disk, and this linearly growing trend was proposed to be removed by an appropriate postprocessing. It turns out that this undesirable behavior was merely a consequence of the numerical implementation (Simpson’s quadrature rule should have been avoided due to the low regularity of the noisy field) and not the asymptotic formula itself, as demonstrated in the present work; the same is also true for higher-order estimates of $m_{3}$. The drawback of finiteness of the measurement region for the application of Helbig’s integral formulas has been recognised a while ago and its influence has been analysed, see, e.g., [17]. One possibility to deal with this was to choose and correct a suitable integration window so that the consistency of Helbig’s integrals is respected (since some integrals of the field data must be identically zero), see [20]. Another approach consists in replacing the missing field data by the field of a fictious dipole whose location and moment are estimated iteratively from appropriate integrals of the available data [1, 9]. A similar iterative strategy [18] is based on the multipolar expansion of order 2 (involving quadrupole moment tensor). In a recent work [16], a multipolar approach of arbitrary order has been used to fit the measured data furnishing the net moment as first coefficients of the multipolar expansion. This shows good results but requires a potentially delicate parameter tunning (such as choice of the origin and the order of the expansion) as a part of the regularisation process. A totally different method which does not theoretically require a large measurement region is that based on a bounded extremal problem for finding best-possible linear estimators [2]. These estimators are auxiliary functions whose integrals against the measured data furnish the components of the net moment. Finding such estimators in a particular functional class requires solving an integro-differential equation which is a numerically laborous task. While formally applicable for any size and position of the measurement region with respect to the magnetic source, a practical sensitivity of that approach to perturbations of the measured data must increase enormously when the measurement area is not sufficiently large or well-located with respect to the source. Moreover, this approach is restricted to planar and regular (square-integrable) sources at a known height (depth).
@@
The present approach is applicable for one-component field measurements. It has high accuracy already for relatively small measurement area due to the higher-order asymptotic estimates (which, unlike Helbig’s integrals, include also estimates for $m_{3}$ from $B_{3}$). The approach is easy to implement numerically due to the explicit nature of the estimates. It is applicable for bounded volumetric magnetic sources which are not necessarily regular (they may be even compactly supported distributions, not just square-integrable functions). Moreover, the approach does not require the knowledge of the depth of the source (as all the estimates are seen to be independent of the height parameter $h$ which must only satisfy a certain a priori assumption).
@@ 2. Main results
Let $\mathcal{E}^{\prime}\left(\mathbb{R}^{3}\right)$ be the space of compactly supported distributions on $\mathbb{R}^{3}$, i.e., the linear functionals on smooth functions $C^{\infty}\left(\mathbb{R}^{3}\right)$, see [25, Sect. 6.1].
Denote $D_{A}:=\left\{\mathbf{x}\in\mathbb{R}^{2}:\,\left|\mathbf{x}\right|<A\right\}$, the disk of radius $A$ centered at the origin $\mathbf{x}=\mathbf{0}$.
@@
Assume that $\vec{\mathcal{M}}\in\left[\mathcal{E}^{\prime}\left(\mathbb{R}^{3}\right)\right]^{3}$ with supp $\text{supp }\vec{\mathcal{M}}\subset Q$ for a bounded set $Q\subset\mathbb{R}^{3}$. Suppose that the values of the vertical component of the magnetic field $B_{3}$ (related to $\vec{\mathcal{M}}$ by means of (1.2)) are known on the horizontal disk $D_{A}\times\left\{x_{3}=h\right\}$ that does not intersect $Q$ and whose radius $A$ is sufficiently large so that the following inequality holds:
(2.1) $\sup_{\vec{t}\in Q,\;\mathbf{x}\in\mathbb{R}^{2}\backslash D_{A}}\left|\frac{t_{1}^{2}+t_{2}^{2}+\left(h-t_{3}\right)^{2}}{x_{1}^{2}+x_{2}^{2}}-2\frac{x_{1}t_{1}+x_{2}t_{2}}{x_{1}^{2}+x_{2}^{2}}\right|<1.$
@@
[…]
@@
$-1<\frac{t_{1}^{2}+t_{2}^{2}+\left(h-t_{3}\right)^{2}}{x_{1}^{2}+x_{2}^{2}}-2\frac{x_{1}t_{1}+x_{2}t_{2}}{x_{1}^{2}+x_{2}^{2}}<1,\hskip 30.00005pt\vec{t}\in Q,\,\,\mathbf{x}\in\mathbb{R}^{2}\backslash D_{A}.$
@@
$\left(x_{1}-t_{1}\right)^{2}+\left(x_{2}-t_{2}\right)^{2}+\left(h-t_{3}\right)^{2}>0,\hskip 30.00005pt\vec{t}\in Q,\,\,\mathbf{x}\in\mathbb{R}^{2}\backslash D_{A},$
@@
$x_{1}^{2}+x_{2}^{2}+2\left(x_{1}t_{1}+x_{2}t_{2}\right)+t_{1}^{2}+t_{2}^{2}>2\left(t_{1}^{2}+t_{2}^{2}\right)+\left(h-t_{3}\right)^{2},\hskip 30.00005pt\vec{t}\in Q,\,\,\mathbf{x}\in\mathbb{R}^{2}\backslash D_{A}.$
By the symmetry of the area $\mathbb{R}^{2}\backslash D_{A}$, we can change the signs in front of $x_{1}$ and $x_{2}$. Recognising the complete square on the left-hand side and dividing over the positive expression from the right-hand side, we arrive exactly at the fraction appearing in (2.10) and it only rests to take the infimum.
Finally, let us point out a simplification occuring in the common setting (e.g., when $Q$ is a minimal rectangular parallelipiped or a ball containing the magnetisation support located under the measurement device). Namely, if $Q$ is a bounded connected set whose lowest points lie at $x_{3}=0$ plane and whose horizontal projection $Q_{12}$ contains the center $\mathbf{x}=\mathbf{0}$ of the measurement disk $D_{A}$, then inequality (2.10) can be ensured by imposing a stricter but geometrically simpler condition
@@
where diam $\text{diam }Q_{12}$ denotes the diameter of $Q_{12}$ and $\text{dist}\left(\partial D_{A},Q_{12}\right)$ is the distance between the boundary of the measurement disk and $Q_{12}$. Indeed, by the above assumption on the location of $Q$ and choice of the origin, we have $t_{1}^{2}+t_{2}^{2}\leq\left(\text{diam }Q_{12}\right)^{2}/4$, $\left(h-t_{3}\right)^{2}\leq h^{2}$ for $\vec{t}\in Q$, and hence the inequality
@@ 3.1. Some notation and preparatory transformations
We also note that the Fourier transform of the two-dimensional Poisson kernel is well-known (see, e.g., [25, Sect. 4.2]), that is, for any $H>0$, we have
[…]

### v5

@@ Abstract.
Inverse magnetisation problem consists in inferring information about a magnetic source from measurements of its magnetic field. Unlike a general magnetisation distribution, the total magnetisation (net moment) of the source is a quantity that theoretically can be uniquely determined from the field. At the same time, it is often the most useful quantity for practical applications (on large and small scales) such as detection of a magnetic anomaly in magnetic prospection problem or finding the overall strength and mean direction of the magnetisation distribution of a magnetised rock sample. It is known that the net moment components can be explicitly estimated using the so-called Helbig’s integrals which involve integration of the magnetic field data on the plane against simple polynomials. Evaluation of these integrals requires knowledge of the magnetic field data on a large region or the use of ad hoc methods to compensate for the lack thereof. In this paper, we derive higher-order analogs of Helbig’s integrals which permit estimation of total magnetisation components in terms of measurement data available on a smaller region. Motivated by a concrete experimental setup for analysing remanent magnetisation of rock samples with a scanning microscope, we also extend Helbig’s integrals to the situation when knowledge of only one field component is necessary. Moreover, apart from derivation of these novel formulas, we rigorously prove their accuracy. The presented approach, based on an appropriate splitting in the Fourier domain and estimates of oscillatory integrals (involving both small and large parameters), elucidates the derivation of asymptotic formulas for the net moment components to an arbitrary order, a possibility that was previously unclear. The obtained results are illustrated numerically and their robustness with respect to the noise is discussed.
@@ 1. Introduction
In the present work, we are concerned with recovery of the total magnetisation (the so-called net moment) of a sample rather than dealing with reconstruction of the magnetisation distribution. While both inverse problems are known to be ill-posed (as an inverse source problem for elliptic partial-differential equations) due to the lack of continuous dependence of their solution on the input data (magnetic field measurements), the problem of reconstruction of the magnetisation distribution additionally lacks uniqueness of the solution in view of presence of invisible (or “silent”) sources, i.e., magnetisations that do not produce magnetic field, see [4]. However, as it was shown in [2] for planar (thin-plate) magnetisation distributions, and as follows from the general Helmholtz decomposition [7], compactly supported invisible sources do not contribute to the net moment. This statement fixes the non-uniqueness issue and makes the problem of net moment recovery a feasible task. In theory, this problem is even solvable in a closed form when measurements are available on the entire plane above the sample. In reality, however, the measurements are very limited and corrupted by the presence of noise which dominates the signal in distant regions. Therefore, we arrive at the problem of estimating the net moment of a sample from a magnetic field component available on some portion of the plane in proximity of the sample. We shall assume that this portion of the plane is sufficiently large, for otherwise the instability of the problem, due to the already mentioned inherent ill-posedness, will be even more severe. On the other hand, this assumption on the large size of the measurement area allows us to obtain explicit ready-to-use formulas.
@@
(1.1) $\Q¯.B_{3}\left(\vec{x}\right)=\dfrac{\partial}{\partial x_{3}}\iiint_{Q}\dfrac{1}{4\pi\left|\vec{x}-\vec{t}\right|}\nabla\cdot\vec{\mathcal{M}}\left(\vec{t}\right)\mathrm{d}^{3}t,\hskip 10.00002pt\vec{x}\in\mathbb{R}^{3}\backslash\overline${Q}.
@@
This is a constant $\mathbb{R}^{3}$ vector equal to the total magnetisation of the sample which can be determined uniquely, unlike the complete magnetisation distribution $\vec{\mathcal{M}}$ which essentially enters the expression for $B_{3}$ through $\nabla\cdot\vec{\mathcal{M}}$ and hence can be changed by any divergence-free vector field without altering values of $B_{3}$, see (1.1) and discussions in [4]. In what follows, we may interchangeably use slightly different names for $\vec{m}$: net moment, net magnetisation, zeroth-order algebraic moment of a magnetisation distribution. When $\vec{\mathcal{M}}$ is not a regular function but a distribution (generalised function), the integral in (1.3) should be understood componentwise as the duality pairing of a compactly supported distribution with the constant function $1$ on $\mathbb{R}^{3}$: $m_{j}=\left\langle\mathcal{M}_{j},1\right\rangle$, $j\in\left\{1,2,3\right\}$.
@@
It should be mentioned that even though recovery of the net moment of a sample, our primary concern, is an important practical problem on its own right, it can also serve, under appropriate assumptions, as a preliminary step for the full magnetisation inversion (i.e., finding a magnetisation distribution $\mathcal$$\vec${M}} that would be consistent with the measured data $B_{3}$ according to (1.2)). Indeed, while it is unrealistic to retrieve three generally independent magnetisation components (functions or, more generally, distributions), $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\mathcal{M}_{3}$, from the partial knowledge of only one function $B_{3}$ (see (1.2)) without additional assumptions, the problem simplifies significantly for a class of samples which are unidirectionally magnetised (i.e., when magnetisation direction does not change throughout the sample but its magnitude does). Since quantifying the net moment of a sample implies a definite magnetisation direction, it is thus an essential element in this complete reconstruction procedure, see [14].
@@
We stress that our analytical expressions for the asymptotic estimates are original, with only partial analogs in other works when translated into the current context. According to [21], for extended finite magnetic sources, net moment estimates in terms of the measured field have first appeared in Helbig’s work [13] and later rediscovered by Clark and Schmidt, see, e.g., [24] in context of magnetic prospection and detection of potential field anomalies, in particular. When truncated to a finite measurement area, these estimates correspond to our first-order asymptotic formulas for tangential components of net moment, but without a mention of the order of the asymptotic convergence, let alone a possibility of obtaining their more precise higher-order counterparts. Note that, in our context, since only $B_{3}$ component of the field is measured, those previous estimates do not provide means to estimate the normal net moment component $m_{3}$. Indeed, an estimate of $m_{3}$ was provided only in terms of either $B_{1}$ or $B_{2}$ components. When measurements of these two field components are not available, one can potentially argue that it is still possible to use the Helbig’s formulas after reconstructing the missing components of the field from the knowledge of $B_{3}$ alone using, for example, [12] and [8, Ch. 12]. However, this is possible only to accomplish in a stable way only if $B_{3}$ is known on the entire plane. In view of this, we believe that an explicit estimate of $m_{3}$ from knowledge of $B_{3}$ alone on a finite region for the first time was proposed only in [22, 5, 3]. It should be noted that results in [3] are also applicable for measurement regions different from disk, namely, for those of rectangular and diamond shapes. We take advantage of this comparison of results to mention and to correct a couple of issues in [5]. First of all, there is a typo in the $m_{3}$ formula obvious when compared to (2.4) or [22, Thm 3.1.1]: in eq. [5, Eq. (5)], the factor $A$ should be in the numerator rather than the denominator. Moreover, the numerical illustration of the $m_{3}$ estimate, in case of noisy data, showed a divergent behaviour with the growth of the measurement disk, and this linearly growing trend was proposed to be removed by an appropriate postprocessing. It turns out that this undesirable behaviour was merely a consequence of the numerical implementation (Simpson’s quadrature rule should have been avoided due to the low regularity of the noisy field) and not the asymptotic formula itself, as demonstrated in the present work; the same is also true for higher-order estimates of $m_{3}$. The drawback of finiteness of the measurement region for the application of Helbig’s integral formulas has been recognised a while ago and its influence has been analysed, see, e.g., [18]. One possibility to deal with this was to choose and correct a suitable integration window so that the consistency of Helbig’s integrals is respected (since some integrals of the field data must be identically zero), see [21]. Another approach consists in replacing the missing field data by the field of a fictitious dipole whose location and moment are estimated iteratively from appropriate integrals of the available data [1, 10]. A similar iterative strategy [19] is based on the multipolar expansion of order 2 (involving quadrupole moment tensor). In a recent work [17], a multipolar approach of arbitrary order has been used to fit the measured data furnishing the net moment as first coefficients of the multipolar expansion. This shows good results but requires a potentially delicate parameter tuning (such as choice of the origin and the order of the expansion) as a part of the regularisation process. A totally different method which does not theoretically require a large measurement region is that based on a bounded extremal problem for finding best-possible linear estimators [2]. These estimators are auxiliary functions whose integrals against the measured data furnish the components of the net moment. Finding such estimators in a particular functional class requires solving an integro-differential equation which is a numerically laborous task. While formally applicable for any size and position of the measurement region with respect to the magnetic source, a practical sensitivity of that approach to perturbations of the measured data must increase enormously when the measurement area is not sufficiently large or well-located with respect to the source. Moreover, this approach is restricted to planar and regular (square-integrable) sources at a known height (depth).
@@
The present approach is applicable for one-component field measurements. It has high accuracy already for relatively small measurement area due to the higher-order asymptotic estimates (which, unlike Helbig’s integrals, include also estimates for $m_{3}$ from $B_{3}$). The approach is easy to implement numerically due to the explicit (closed-form) nature of the estimates. It is applicable for bounded volumetric (and planar, in particular case) magnetic sources which are not necessarily regular (they may be even compactly supported distributions, not just square-integrable functions). Moreover, the approach does not require the knowledge of the depth of the source (as all the estimates are seen to be independent of the height parameter $h$ which must only satisfy a certain a priori assumption). It is worth stressing that applicability of the approach to volumetric samples is not a purely academic generalisation of that for planar (thin-slab) magnetic distributions: it is important for practical situations such as QDM microscopy (when the lift-off distance is too small, contributions from deeper sources are not negligible) but also in macroscopic contexts such as the already mentioned magnetic prospection problem.
@@ 2. Main results
Let $\mathcal{E}^{\prime}\left(\mathbb{R}^{3}\right)$ be the space of compactly supported distributions on $\mathbb{R}^{3}$, i.e., the linear functionals on smooth functions $C^{\infty}\left(\mathbb{R}^{3}\right)$, see [26, Sect. 6.1].
Denote $D_{A}:=\left\{\mathbf{x}\in\mathbb{R}^{2}:\,\left|\mathbf{x}\right|<A\right\}$, the disk of radius $A$ centreed at the origin $\mathbf{x}=\mathbf{0}$.
@@
Assume that $\vec{\mathcal{M}}\in\left[\mathcal{E}^{\prime}\left(\mathbb{R}^{3}\right)\right]^{3}$ with $\text{supp }\vec{\mathcal{M}}\subset Q$ for a bounded set $Q\subset\mathbb{R}^{3}$. Suppose that the values of the vertical component of the magnetic field $B_{3}$ (related to $\vec{\mathcal{M}}$ by means of (1.2)) are known on the horizontal disk $D_{A}\times\left\{x_{3}=h\right\}$ that does not intersect $Q$ and whose radius $A$ is sufficiently large so that the following inequality holds:
(2.1) $sup_{\vec{t}\in Q,\;\mathbf{x}\in\mathbb{R}^{2}\backslash D_{A}}\left|\frac{t_{1}^{2}+t_{2}^{2}+\left(h-t_{3}\right)^{2}}{x_{1}^{2}+x_{2}^{2}}-2\frac{x_{1}t_{1}+x_{2}t_{2}}{x_{1}^{2}+x_{2}^{2}}\right|<1.$
@@
(2.6) $m_{3}=\frac{A}{4}\iint_{D_{A}}\left[5+40\left(\frac{x_{j}}{A}\right)^{4}-128\left(\frac{x_{j}}{A}\right)^{6}\right]B_{3}\left(\mathbf{x},h\right)\mathrm{d}^{2}x+\mathcal{O}\left(\frac{1}{A^{3}}\right),\hskip 10.00002pt\hskip 10.00002ptj\in\left\{1,2\right\}.$
@@
[…]
@@
[…]
@@
[…]
Finally, let us point out a simplification occurring in the common setting (e.g., when $Q$ is a minimal rectangular parallelepiped or a ball containing the magnetisation support located under the measurement device). Namely, if $Q$ is a bounded connected set whose lowest points lie at $x_{3}=0$ plane and whose horizontal projection $Q_{12}$ contains the centre $\mathbf{x}=\mathbf{0}$ of the measurement disk $D_{A}$, then inequality (2.10) can be ensured by imposing a stricter but geometrically simpler condition
@@
where $\text{diam }Q_{12}$ denotes the diameter of $Q_{12}$ and $\text{dist}\left(\partial D_{A},Q_{12}\right)$ is the distance between the boundary $\partial D_{A}$ of the measurement disk and $Q_{12}$. Indeed, by the above assumption on the location of $Q$ and choice of the origin, we have $t_{1}^{2}+t_{2}^{2}\leq\left(\text{diam }Q_{12}\right)^{2}/4$, $\left(h-t_{3}\right)^{2}\leq h^{2}$ for $\vec{t}\in Q$, and hence the inequality
@@
The derivation and the proof of the main estimates is lengthy and consists of a combination of simple techniques such as Fourier transforms, series expansions (at the origin and at infinity) and appropriate integral splitting. One of the main difficulties that some integrals involve both large and small parameters and hence obtaining uniform estimates by standard means of analysis of oscillatory and decaying integrals fails. Instead, we are able to benefit from circular geometry of the measurement area by identifying integral representations of known special (Bessel type) functions and rely on their differential and integral identities and asymptotics [20, Sect. 10–11] to conclude with the desired estimates.
@@ 3.1. Some notation and preparatory transformations
We also note that the Fourier transform of the two-dimensional Poisson kernel is well-known (see, e.g., [26, Sect. 4.2]), that is, for any $H>0$, we have
(3.4) $\mathcal{F}\left[\frac{H}{2\pi\left(\left|\mathbf{x}\right|^{2}+H^{2}\right)^{3/2}}\right]\left(\mathbf{k}\right)=e^{-2\pi H\left|\mathbf{k}\right|},\hskip 10.00002pt\mathbf{k}\in\mathbb{R}^{2}.$